OpenStudy (anonymous):

someone tell me how to answer (X+27)/5 = 2x in full?

OpenStudy (amistre64):

multiply all by 5; -27 from each side; -10x from eaqh side; and divide it all by -9

OpenStudy (amistre64):

if i see it right, i get 3

OpenStudy (anonymous):

confused ^_^

OpenStudy (amistre64):

with which step?

OpenStudy (anonymous):

so it would be (X+27) x 5 = 2x x 5?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

but how could you times them each by 5? :s

OpenStudy (amistre64):

but you forgot your denominator of 5 that cancels the 5 on top when you do that :)

OpenStudy (amistre64):

let me demonstrate how equeations work ... does 15 = 15?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

if we subtract 3 from both sides; does it still equal each other?

OpenStudy (anonymous):

yup

OpenStudy (amistre64):

if we divide both sides by 300; do they still equal each other?

OpenStudy (anonymous):

yup

OpenStudy (amistre64):

does it matter what we do to them as long as we do it to both of them in the same manner; will they always equal the same amount?

OpenStudy (anonymous):

yeah they will always be the same, i just can't figure out how to times the (x+27) by 5 and the 2x by 5 will = 10x right?

OpenStudy (amistre64):

2x x 5 = 10x yes; (x+27) ------ x 5 = (x+27) 5 soo: x+27 = 10x

OpenStudy (anonymous):

ah right i get you

OpenStudy (anonymous):

so then 27 = 11x?

OpenStudy (amistre64):

now -27 and -10x and divide it all by -9 and get x=3 lol

OpenStudy (anonymous):

how do you get to the -27 and -10x?

OpenStudy (amistre64):

your way was good; just a little off -x from each side to get : 27 = 9x

OpenStudy (amistre64):

x-x+27 = 10-x 27 = 9x

OpenStudy (amistre64):

and thats spose to be a 10x-x :) typoed meself

OpenStudy (anonymous):

right so what i have atm is, x+27 ----- = 2x 5 x+27 ----- x5 = 2x x5 5 = x+27 = 10x = x + 27 - x = 10x - x = 27 = 9x

OpenStudy (amistre64):

yesss, thats correct; now divide each side by 9 to get that value for x :)

OpenStudy (anonymous):

where did you get the divide 9 from though?

OpenStudy (anonymous):

the 9x?

OpenStudy (amistre64):

9x has baggage ; its that 9 stuck to it by multiplication; to undo the multiplication; divide

OpenStudy (amistre64):

9x 9 -- = -- x = 1x = x 9 9

OpenStudy (anonymous):

so then final answer is x = 3?

OpenStudy (amistre64):

well; with the x all alone, thats what we want

OpenStudy (anonymous):

i think i have it :P, thanks so much for that, are you busy as i have a similar one that needs answering should just need tweeking from the previous answer (i think)

OpenStudy (anonymous):

which is (4x+6)/ 3 = 10

OpenStudy (anonymous):

so it would be x3 each side?

OpenStudy (anonymous):

4x + 6 = 30, then i'm correct, 4x = 30 - 6 which = 24, and then 4x = 24 /4 each side, x = 6?

OpenStudy (amistre64):

thats good; always dbl chk the answer by plugging it back into the original and see if it makes a true statement

OpenStudy (amistre64):

(4(6)+6)/3 = 10 (24+6)/3 = 10 30/3 = 10 10=10